# Boosting Visual-Language Models by Exploiting Hard Pairs

### v2

@@ Abstract
Contrastive Language-Image Pre-training (CLIP) has become the standard for learning cross-modal representations between images and text. Efforts to improve its capabilities typically demand the collection of additional data and retraining with new loss functions. While effective, the added requirements limit their practical use due to the increased resource and time investments needed. In this work, we present Helip, a cost-effective strategy tailored to enhance the performance of existing CLIP models without the need for training a model from scratch or collecting additional data. Our method allows for effortless integration with existing models’ training pipelines, providing an instant boost by training them with selected challenging text-image pairs from their original training datasets. Helip treats each text-image pair as a single point in the joint vision-language space, identifying those in close proximity as hard pairs. By incorporating the challenging data, pre-trained CLIP models are refined using both the traditional contrastive loss and the newly introduced hard negative margin loss, ensuring the challenging data is fully utilized. On comprehensive benchmarks, Helip consistently boosts existing models to achieve leading performance. In particular, it improves the zero-shot classification accuracy on ImageNet for SLIP models pre-trained on CC3M, CC12M and YFCC15M datasets. The improvements are $3.05\%$, $4.47\%$, and $10.1\%$ respectively, achieved within two epochs of training. In addition, across fine-grained classification datasets, Helip improves the zero-shot performance of pre-trained CLIP and SLIP by an average of $8.4\%$ and $18.6\%$, and their linear probe performance by an average of $9.5\%$ and $3.0\%$. The code is publicly available at https://github.com/haonan3/HELIP.
@@ 1 Introduction
Contrastive Language-Image Pretraining (CLIP) (Radford et al., 2021) is quickly becoming the standard for foundation models (Awais et al., 2023) due to its effectiveness for a variety of vision-language tasks without task-specific finetuning (Li et al., 2021; Baldrati et al., 2022). However, web-crawled image-text pairs used for the CLIP model pretraining are often loosely connected, resulting in multiple plausible matches beyond the assigned ones (Wu et al., 2022). Several methods have been presented to improve CLIP models by investigating appropriate matches and utilizing widespread supervision among image-text pairs for model training (Li et al., 2022a, 2021; Mu et al., 2022; Radenovic et al., 2023).
Efforts to improve contrastive language-image pretraining models have primarily taken two directions: (1) the addition of objectives to improve the efficacy of supervision (Li et al., 2022a; Mu et al., 2022); and (2) the employment of intra- and inter-modality similarity to select and retrain models using data deemed challenging at the sample level (Li et al., 2021; Radenovic et al., 2023). However, those approaches inevitably require retraining, and those identified as challenging data struggle to bring benefits to model performance. This challenge is partly due to their reliance on finding challenging data within a single batch during training, where truly beneficial challenging data is rare. Additionally, CLIP models’ contrastive loss is not optimally configured to exploit the nuances of difficult data. These limitations significantly restrict the practical application of these enhancements, especially considering the substantial investments already made in pretraining numerous CLIP models (Li et al., 2022a; Mu et al., 2022). This aspect underscores the need for efficient enhancement strategies that do not rely on additional data collection to improve existing pretrained models.
@@
Empirical evidence shows that Helip improves the performance of existing CLIP models, including pre-trained CLIP, SLIP, and DECLIP, across a variety of benchmarks, such as zero-shot classification, text-image retrieval, and fine-grained linear probing. For zero-shot classification on ImageNet, CIFAR-10, and CIFAR-100, Helip consistently boosts the performance of all six pre-trained models. Particularly, using Helip to boost SLIP models pre-trained on CC3M, CC12M, and YFCC15M results in ImageNet zero-shot accuracy gains of $3.05\%$, $4.47\%$, and $10.14\%$, respectively. Further, on seven fine-grained image classification datasets, those pre-trained models achieve better zero-shot and linear probe performance with Helip. Specifically, the average zero-shot accuracy of CC3M pre-trained CLIP and SLIP are improved by $8.4\%$ and $18.6\%$. The average linear probe accuracy of CC3M pre-trained CLIP and SLIP are improved by $9.5\%$ and $3.0\%$ respectively. Additionally, the performance gain is also valid in terms of zero-shot retrieval, with $1.1$ of R@1 on Flickr30K, and $2.2$ of R@1 on COCO for SLIP with Helip. Our contributions could be summarized as:
  • o our best knowledge, our method, Helip  stands out as the first method aimed at improving existing CLIP models in a cost-effective and easily integrable way.
  • We introduce the hard pair mining strategy to select challenging data, accompanied by the development of hard negative margin loss. This combination ensures the effective identification and utilization of challenging data, improving the CLIP models.
  • Empirical evaluations across zero-shot classification, image-text retrieval, and linear probe benchmarks, consistently show Helip’s ability to substantially boost the performance of existing CLIP models, underlining its effectiveness and practicality in real-world applications.
@@ 2 Related work
Vision-Language pre-training. Vision Language Pretraining (VLP) is a technique that leverages large-scale image-text datasets to learn a strong joint representation between the two modalities that can be transferred to various downstream vision-language tasks. VLP models can be generally divided into single-stream models and dual-stream models. Dual-stream models  (Jia et al., 2021; Li et al., 2022b; Mu et al., 2022; Radford et al., 2021; Yao et al., 2022) typically consist of two separate encoders for image and text respectively and perform cross-modality interactions on the top, are becoming more and more popular because of its flexibility of transferring pre-trained knowledge to downstream tasks. CLIP  (Radford et al., 2021), uses a simple contrastive objective to learn visual features from natural language supervision and achieves remarkable zero-shot recognition performance using 400M web-crawled image-text pairs. Recent works boot the performance of CLIP by applying self-supervision within visual modal  (Mu et al., 2022), additional nearest neighbor supervision (Li et al., 2022b). These methods are actually doing data augmentations to increase data efficiency and thus bring additional computational costs.
Contrastive learning with hard negative samples. Contrastive learning learns a representation of input data that maps semantically comparable examples close together and semantically dissimilar examples far apart (Chen et al., 2020a, b; Wang & Isola, 2020). Recent works include hard negative samples into the loss function and achieve better empirical performance (Cai et al., 2020; Huynh et al., 2022; Kalantidis et al., 2020; Li et al., 2021; Radenovic et al., 2023; Robinson et al., 2021; Shah et al., 2022). For Language-image contrastive learning, current approaches (Li et al., 2021; Radenovic et al., 2023) mine multimodal hard negative examples using intra/inter-modality similarity.  Li et al. (2021) choose in-batch hard negative samples with image-text contrastive loss. Hard negative noise contrastive multimodal alignment loss by  Radenovic et al. (2023) up-weights the loss term for in-batch hard samples. For previous intra/inter-modality hard sample mining methods, two text-image pairs are considered as hard samples, if the cosine similarity between visual/textual features is high (Li et al., 2021; Radenovic et al., 2023). However, due to the nature of loose assignment for web-crawled image-caption data, a high similarity indicated by intra/inter-modality doesn’t indicate that the two pairs are difficult to tell apart. Contrary to prior works, we design a hard sample mining method to discover similar pairs defined in joint vision-language space and efficiently select samples challenging enough to improve learning.
@@ 3 Hard pairs for visual-language models
In this section, we first define the notations and revisit CLIP for zero-shot recognition in the preliminary section. Next, we introduce the Hard Pairs Mining strategy (HPM), and the associated Hard Negative Margin Loss (HNML), specifically designed to leverage hard pairs identified by HPM.
@@ 3.1 Preliminaries
$\ell_{CLIP}=-\frac{1}{|B|}\sum_{i\in B}\log\frac{\exp\left(sim(I_{i},T_{i})/\sigma\right)}{\sum_{j\in B}\exp\left(sim(I_{i},T_{j})/\sigma\right)},$ (1)
@@ 3.2 HPM: hard pair mining
Selecting hard pairs by maximizing pair agreement. To identify nearby pairs, we introduce the idea of text-image pair agreement maximization. This can be viewed as a proxy task for selecting hard pairs. To illustrate the rationale for using text-image pair agreement as a proxy for selecting hard pairs, we return to the principle obtained from traditional machine learning methods: the prediction of a model on a test sample is substantially influenced by samples in the training dataset that are similar to the test one. For example, the K-Nearest Neighbors (KNN) algorithm classifies a new instance using the K-closest training examples. The linear regression model predicts the output of a test sample using the weighted sum of the training samples, with higher weights given to samples that are more similar to the test sample. Recent empirical and theoretical studies on model memorization and generalization (Chen et al., 2009; Zhang et al., 2021; Stephenson et al., 2021; Brown et al., 2021) also provide support for this. Intuitively, if a pair agreement prediction model trained on a set of pairs predicts a specific target pair as having a high probability of being a matching pair, the target pair is likely to be similar to the matching pairs on which the model was trained. The challenge of selecting hard pairs is transformed into an optimization task centered on the text-image pair agreement, which is formally represented as:
@@
where $P_{\mathcal{M}}(z_{i}|\mathcal{H}_{i})$ denotes the prediction of a pair agreement model, $\mathcal{M}$, for the pair $z_{i}$ based on a pair set $\mathcal{H}_{i}$. This set is a subset of ${\mathcal{D}}_{i}$. In this framework, the goal of selecting hard pair is transformed into identifying a training set $\mathcal{H}_{i}$ such that the model $\mathcal{M}$ predicting the target pair as a matching pair.
@@
Designing a suitable pair agreement prediction model for this proxy task is a nontrivial endeavor because the model needs to not only predict the pair matching probability but also allow the optimization of the training set, as indicated in Equation 3. Consequently, a conventional deep neural network design becomes unviable due to the impracticality of retraining across all possible sets $\mathcal{H}_{i}$ from ${\mathcal{D}}_{i}$. Taking inspiration from recent work (Norelli et al., 2022), we propose a data-centric design for the agreement prediction model $\mathcal{M}$. As illustrated in Figure 1, the model leverages two pretrained single-modal encoders, i.e., $f_{image}$ and $f_{text}$, to align representations of images and texts in a unified Vision-Language space. Specifically, the model encodes the target pair $z_{i}$ into $(I_{i},T_{i})$ using these single-modal encoders. For the visual modality, we determine a similarity vector between the target pair $z_{i}$ and the dataset ${\mathcal{D}}_{i}$. The similarity vector is defined as $\vec{S}^{I}(x^{I}_{i},{\mathcal{D}}_{i})=[\dots,sim(I_{i},I_{j}),\dots]^{\top}\in\mathbb{R}^{N-1}$. Here $I_{j}=f_{image}(x^{I}_{j})$ with $(x^{I}_{j},x^{T}_{j})$ being an element of ${\mathcal{D}}_{i}$, and function $sim(\cdot,\cdot)$ denotes the cosine similarity. To counteract noise, values in the vector $\vec{S}^{I}(x^{I}_{i},{\mathcal{D}}_{i})$ are set to zero if $sim(I_{i},I_{j})<\tau$. This cleaned-up vector is represented as $\widetilde{S}^{I}$. The procedure for the textual modality is analogous, producing a vector denoted as $\widetilde{S}^{T}$. Note, the representations in this shared space are intuitively interpretable: each dimension corresponding to the visual/textual similarity of the input to a unique pair in the multimodal dataset. This interpretable characteristic enables us to directly optimize the supporting set to maximize the pair matching probability:
$\mathcal{H}^{\star}_{i}=Argmax_{|\mathcal{H}_{i}|=k}\widetilde{S}^{I}(x^{I}_{i},\mathcal{H}_{i})^{\top}\widetilde{S}^{T}(x^{T}_{i},\mathcal{H}_{i}),$ (4)
@@
where the $\mathcal{H}^{\star}_{i}$ is the hard pair set and $k\in\mathbb{R}^{+}$ is the number of selected pairs which is much less than $|\mathcal{D}|$. The previous problem can be efficiently solved by greedily choosing dimensions that maximize the inner product. Due to the interpretable property, the selected dimensions correspond to the desired pairs.
@@
Mitigation of noisy data impact. The prior method assumes the target pair $z_{i}$ to be a suitable matching pair. However, in inherently noisy datasets, such as web-crawled ones like LAION (Schuhmann et al., 2022), mismatched pairs might be present. The potential negative effects of hard pairs generated by these mismatched pairs necessitate the development of a strategy for identifying and eliminating them. We create a pair removal strategy based on the availability of hard pairs: A target pair $z_{i}$ is deemed as unsuitable and thus removed, if there is a non-empty subset $\mathcal{H}^{sub}_{i}\subseteq\mathcal{H}^{\star}_{i}$ with $|\mathcal{H}^{sub}_{i}|>0$, such that $\widetilde{S}^{I}(x^{I}_{i},\mathcal{H}^{sub}_{i})^{\top}\widetilde{S}^{T}(x^{T}_{i},\mathcal{H}^{sub}_{i})=0$. Intuitively, this equation indicates that the number of matching pairs that support $z_{i}$ is less than $k$. If there is no small subset of size $k$ to support $z_{i}$ as a matching pair in dataset $\mathcal{D}\setminus z_{i}$, it indicates that the target pair is an outlier, possibly due to a mismatch. Such outliers can degrade dataset quality, so they are removed to ensure the reliability of hard data.
@@
$\mathcal{H}^{\star}_{i}\approx Argmax_{|\mathcal{H}|=k}\widetilde{S}^{I}(x^{I}_{i},\mathcal{H}_{i})^{\top}\widetilde{S}^{T}(x^{T}_{i},\mathcal{H}_{i}),$ (5)
where $\mathcal{H}_{i}\subseteq\overline{\mathcal{D}}_{i}$ and $|\overline{\mathcal{D}}_{i}|=C$ is sampled uniformly from set ${\mathcal{D}}_{i}$. In this equation, it’s noteworthy that the selection of value $C$ is solely based on the number of hard pairs $k$, instead of the size of ${\mathcal{D}}_{i}$. Consequently, this optimization reduces the time complexity of FastHPM to $\mathcal{O}(N)$. The detailed procedure of the hard pair mining algorithm is consolidated and presented in Appendix A.

### v3

# Getting More Juice Out of Your Data: Hard Pair Refinement Enhances Visual-Language Models Without Extra Data

@@ Abstract
Contrastive Language-Image Pre-training (CLIP) has become the standard for cross-modal image-text representation learning. Improving CLIP typically requires additional data and retraining with new loss functions, but these demands raise resource and time costs, limiting practical use. In this work, we introduce Helip, a cost-effective strategy that improves CLIP models by exploiting challenging text-image pairs within existing datasets in continuous training. This eliminates the need for additional data or extensive retraining. Moreover, Helip integrates effortlessly into current training pipelines with minimal code modifications, allowing for quick and seamless implementation. On comprehensive benchmarks, Helip consistently boosts existing models. In particular, within just two epochs of training, it improves zero-shot classification accuracy on ImageNet for SLIP models pre-trained on CC3M, CC12M, and YFCC15M datasets by $3.05\%$, $4.47\%$, and $10.1\%$ , respectively. In addition, on fine-grained classification datasets, Helip improves the zero-shot performance of CLIP and SLIP by an average of $8.4\%$ and $18.6\%$, and their linear probe performance by an average of $9.5\%$ and $3.0\%$. The code is publicly available at: https://github.com/haonan3/HELIP-NAACL-2025.git.
Getting More Juice Out of Your Data: Hard Pair Refinement Enhances Visual-Language Models Without Extra Data
Haonan Wang1thanks: Equally contributed to this work, Corresponding author.    §Affiliated with Department of Systems Engineering and Engineering Management, and Shun Hing Institute of Advanced Engineering, The Chinese University of Hong Kong, Hong Kong, Minbin Huang2∗§, Runhui Huang3, Lanqing Hong4$\dagger$, Hang Xu4, Tianyang Hu1, Xiaodan Liang3, Zhenguo Li3, Hong Cheng2§, Kenji Kawaguchi1 1National University of Singapore  2The Chinese University of Hong Kong 3Sun Yat-sen University  4Huawei Noah’s Ark Lab
@@ 1 Introduction
Contrastive Language-Image Pretraining (CLIP) (Radford et al., 2021) is quickly becoming the standard for foundation models (Awais et al., 2023) due to its effectiveness for a variety of vision-language tasks without task-specific finetuning (Li et al., 2021; Baldrati et al., 2022). However, web-crawled image-text pairs used for the CLIP model pretraining are often loosely connected, resulting in multiple plausible matches beyond the assigned ones (Wu et al., 2022). Several methods have been presented to improve CLIP models by investigating appropriate matches and utilizing widespread supervision among image-text pairs for training (Li et al., 2022a, 2021; Mu et al., 2022; Radenovic et al., 2023).
Efforts to improve contrastive language-image pretraining models have primarily taken two directions: (1) the addition of objectives to improve the efficacy of supervision (Li et al., 2022a; Mu et al., 2022); and (2) the employment of intra- and inter-modality similarity to select and retrain models using data deemed challenging at the sample level (Li et al., 2021; Radenovic et al., 2023). However, those approaches inevitably require retraining, and those identified as challenging data struggle to bring benefits to model performance. This challenge is partly due to their reliance on finding challenging data within a single batch during training, where truly beneficial challenging data is rare. And, CLIP models’ original contrastive loss is not optimally configured to exploit the nuances of difficult data. These limitations restrict the practical application of these methods, especially considering the substantial investments already made in pretraining numerous CLIP models (Li et al., 2022a; Mu et al., 2022); retraining for minimal gains is inefficient. This aspect underscores the need for efficient enhancement strategies that do not rely on additional data collection to improve existing pretrained models.
@@
Empirical evidence shows that Helip improves the performance of existing CLIP models, including pre-trained CLIP, SLIP, and DECLIP, across a variety of benchmarks, such as zero-shot classification, text-image retrieval, and fine-grained linear probing. For zero-shot classification on ImageNet, CIFAR-10, and CIFAR-100, Helip consistently boosts the performance of all six pre-trained models. Particularly, using Helip to boost SLIP models pre-trained on CC3M, CC12M, and YFCC15M results in ImageNet zero-shot accuracy gains of $3.05\%$, $4.47\%$, and $10.14\%$, respectively. Further, on seven fine-grained image classification datasets, those pre-trained models achieve better zero-shot and linear probe performance with Helip. Specifically, the average zero-shot accuracy of CC3M pre-trained CLIP and SLIP are improved by $8.4\%$ and $18.6\%$. The average linear probe accuracy of CC3M pre-trained CLIP and SLIP are improved by $9.5\%$ and $3.0\%$ respectively. Additionally, the performance gain is also valid in terms of zero-shot retrieval, with $1.1$ of R@1 on Flickr30K, and $2.2$ of R@1 on COCO for SLIP with Helip.
@@ 2 Related Work
Vision-Language Pre-training. Vision-Language Pretraining (VLP) leverages large-scale image-text datasets to learn joint representations transferable to downstream tasks. VLP models are typically classified into single-stream and dual-stream architectures. Single-stream models concatenate text and visual features processed by a single transformer (Li et al., 2019; Chen et al., 2022; Zhang et al., 2020). Dual-stream models use separate encoders for image and text, performing cross-modal interactions at a higher level (Radford et al., 2021; Jia et al., 2021; Li et al., 2022b; Mu et al., 2022; Yao et al., 2022). CLIP (Radford et al., 2021), a dual-stream model, employs contrastive learning with 400M web-crawled image-text pairs to achieve remarkable zero-shot recognition performance. Recent works enhance CLIP’s performance by applying self-supervision within the visual modality (Mu et al., 2022) or incorporating nearest neighbor supervision (Li et al., 2022b). While these methods improve performance, they introduce additional computational costs due to data augmentations.
Contrastive Learning with Hard Negative Samples. Contrastive learning aims to learn representations by bringing similar examples closer and pushing dissimilar ones apart (Chen et al., 2020a, b; Wang and Isola, 2020). Incorporating hard negative samples into the loss function has been shown to improve performance (Cai et al., 2020; Huynh et al., 2022; Kalantidis et al., 2020; Li et al., 2021; Radenovic et al., 2023; Robinson et al., 2021; Shah et al., 2022). In language-image contrastive learning, approaches like Li et al. (2021) and Radenovic et al. (2023) mine hard negatives using intra- or inter-modality similarity, selecting samples with high cosine similarity in visual or textual features. However, due to the loose alignment in web-crawled data, high similarity in these features doesn’t necessarily indicate that pairs are difficult to distinguish. In contrast, we propose a hard sample mining method that discovers similar pairs in the joint vision-language space, efficiently selecting truly challenging samples to improve learning.
@@ 3 Hard Pairs for Visual-Language Models
In this section, we first define the notations and revisit CLIP for zero-shot recognition in the preliminary section. Next, we introduce the Hard Pairs Mining strategy (HPM), and the associated Hard Negative Margin Loss (HNML), designed to efficiently exploit hard pairs.
@@ 3.1 Preliminaries
$\ell_{\text{CLIP}}=-\frac{1}{|B|}\sum_{i\in B}\log\frac{\exp\left(sim(I_{i},T_{i})/\sigma\right)}{\sum_{j\in B}\exp\left(sim(I_{i},T_{j})/\sigma\right)},$ (1)
@@ 3.2 HPM: Hard Pair Mining
To identify nearby pairs, we introduce the idea of text-image pair agreement maximization. This can be viewed as a proxy task for selecting hard pairs. To illustrate the rationale for using text-image pair agreement as a proxy for selecting hard pairs, we return to the principle obtained from traditional machine learning methods: the prediction of a model on a test sample is substantially influenced by samples in the training dataset that are similar to the test one. For example, the K-Nearest Neighbors (KNN) algorithm classifies a new instance using the K-closest training examples. The linear regression model predicts the output of a test sample using the weighted sum of the training samples, with higher weights given to samples that are more similar to the test sample. Recent empirical and theoretical studies on model memorization and generalization (Chen et al., 2009; Zhang et al., 2021; Stephenson et al., 2021; Brown et al., 2021) also provide support for this. Intuitively, if a pair agreement prediction model trained on a set of pairs predicts a specific target pair as having a high probability of being a matching pair, the target pair is likely to be similar to the matching pairs on which the model was trained. The challenge of selecting hard pairs is transformed into an optimization task centered on the text-image pair agreement, which is formally represented as:
@@
where $P_{\mathcal{M}}(z_{i}|\mathcal{H}_{i})$ denotes the prediction of a pair agreement model, $\mathcal{M}$, for the pair $z_{i}$ based on a pair set $\mathcal{H}_{i}$. This set is a subset of ${\mathcal{D}}_{i}$. In this framework, the goal of selecting a hard pair is transformed into identifying a training set $\mathcal{H}_{i}$ such that the model $\mathcal{M}$ predicts the target pair as a matching pair.
@@
Designing a suitable pair agreement prediction model for this proxy task is a nontrivial endeavor because the model needs to not only predict the pair matching probability but also allow the optimization of the training set, as indicated in Equation 3. Consequently, a conventional deep neural network design becomes unviable due to the impracticality of retraining across all possible sets $\mathcal{H}_{i}$ from ${\mathcal{D}}_{i}$. Taking inspiration from recent work (Norelli et al., 2022), we propose a data-centric design for the agreement prediction model $\mathcal{M}$. As illustrated in Figure 1, the model leverages two pretrained single-modal encoders, i.e., $f_{\text{image}}$ and $f_{\text{text}}$, to align representations of images and texts in a unified Vision-Language space. Specifically, the model encodes the target pair $z_{i}$ into $(I_{i},T_{i})$ using these single-modal encoders. For the visual modality, we determine a similarity vector between the target pair $z_{i}$ and the dataset ${\mathcal{D}}_{i}$. The similarity vector is defined as $\vec{S}^{I}(x^{I}_{i},{\mathcal{D}}_{i})=[\dots,sim(I_{i},I_{j}),\dots]^{\top}\in\mathbb{R}^{N-1}$. Here $I_{j}=f_{\text{image}}(x^{I}_{j})$ with $(x^{I}_{j},x^{T}_{j})$ being an element of ${\mathcal{D}}_{i}$, and function $sim(\cdot,\cdot)$ denotes the cosine similarity. To counteract noise, values in the vector $\vec{S}^{I}(x^{I}_{i},{\mathcal{D}}_{i})$ are set to zero if $sim(I_{i},I_{j})<\tau$. This cleaned-up vector is represented as $\widetilde{S}^{I}$. The procedure for the textual modality is analogous, producing a vector denoted as $\widetilde{S}^{T}$. Note, the representations in this shared space are intuitively interpretable: each dimension corresponding to the visual/textual similarity of the input to a unique pair in the multimodal dataset. This interpretable characteristic enables us to directly optimize the supporting set to maximize the pair matching probability:
$\mathcal{H}^{\star}_{i}=\operatorname*{arg\,max}_{|\mathcal{H}_{i}|=k}\widetilde{S}^{I}(x^{I}_{i},\mathcal{H}_{i})^{\top}\widetilde{S}^{T}(x^{T}_{i},\mathcal{H}_{i}),$ (4)
@@
where the $\mathcal{H}^{\star}_{i}$ is the hard pair set and $k\in\mathbb{R}^{+}$ is the number of selected pairs which is much less than $|\mathcal{D}|$. The previous problem can be efficiently solved by greedily choosing dimensions that maximize the inner product. Due to the interpretable property, the selected dimensions are corresponding to the desired pairs.
@@
Mitigation of Noisy Data Impact. The prior method assumes the target pair $z_{i}$ to be a suitable matching pair. However, in inherently noisy datasets, such as web-crawled ones like LAION (Schuhmann et al., 2022), mismatched pairs might be present. The potential negative effects of hard pairs generated by these mismatched pairs necessitate the development of a strategy for identifying and eliminating them. We create a pair removal strategy based on the availability of hard pairs: A target pair $z_{i}$ is deemed as unsuitable and thus removed, if there is a non-empty subset of the mined hard pair set, $\mathcal{H}^{sub}_{i}\subseteq\mathcal{H}^{\star}_{i}$ with $|\mathcal{H}^{sub}_{i}|>0$, such that $\widetilde{S}^{I}(x^{I}_{i},\mathcal{H}^{sub}_{i})^{\top}\widetilde{S}^{T}(x^{T}_{i},\mathcal{H}^{sub}_{i})=0$.
Intuitively, this equation suggests that the number of entries positively supporting the target pair $z_{i}$ as a matching pair is fewer than $k$. To illustrate how this concept can aid in cleaning noisy data, consider the following example: Suppose the target pair consists of a “cat” image but a “dog” caption (clearly it is a mismatch). For it to be considered a correct match, numerous pairings with same erroneous pattern (i.e., “cat” images paired with “dog” captions) would need to exist in the dataset. By assuming a certain error types are fewer than $k$ throughout the dataset, if no subset of size $k$ within the dataset $\mathcal{D}\setminus z_{i}$ supports $z_{i}$ as a matching pair, this signals that the target pair is an outlier, likely due to a labeling error or mismatch. Such outliers can degrade dataset quality, so they are removed to ensure the reliability of hard data.
@@
$\mathcal{H}^{\star}_{i}\approx\operatorname*{arg\,max}_{|\mathcal{H}|=k}\widetilde{S}^{I}(x^{I}_{i},\mathcal{H}_{i})^{\top}\widetilde{S}^{T}(x^{T}_{i},\mathcal{H}_{i}),$ (5)
where $\mathcal{H}_{i}\subseteq\overline{\mathcal{D}}_{i}$ and $|\overline{\mathcal{D}}_{i}|=C$ is sampled uniformly from set ${\mathcal{D}}_{i}$. In this equation, it’s noteworthy that the selection of value $C$ is solely based on the number of hard pairs $k$, instead of the size of ${\mathcal{D}}_{i}$. Consequently, this optimization reduces the time complexity of FastHPM to $\mathcal{O}(N)$. The detailed procedure of the hard pair mining algorithm is presented in Appendix A.